# CAPRI-FAIR: Integration of Multi-sided Fairness in Contextual POI Recommendation Framework

### v1

@@ Abstract.
Point-of-interest (POI) recommendation, a form of context-aware recommendation, takes into account spatio-temporal constraints and contexts like distance, peak business hours, and previous user check-ins. Given the ability of these kinds of systems to influence not just the consumer’s travel experience, but also the POI’s business, it is important to consider fairness from multiple perspectives. Unfortunately, these systems tend to provide less accurate recommendations to inactive users, and less exposure to unpopular POIs. The goal of this paper is to develop a post-filter methodology that incorporates provider and consumer fairness factors into pre-existing recommendation models, to satisfy fairness metrics like item exposure, and performance metrics like precision and distance, making the system more sustainable to both consumers and providers. Experiments have shown that using a linear scoring model for provider fairness in re-scoring recommended items yields the best tradeoff between performance and long-tail exposure, in some cases without a significant decrease in precision. When attempting to address consumer fairness by recommending more popular POIs to inactive users, the result was an increase in precision for only some recommendation models and datasets. Finally, when considering the tradeoff between both parameters, the combinations that reached the Pareto front of consumer and provider fairness, unfortunately, achieved the lowest precision values. We find that the nature of this tradeoff depends heavily on the model and the dataset.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: -; -; -booktitle: 18th ACM Conference on Recommender Systemsccs: Information systems Recommender systemsccs: Information systems Combination, fusion and federated search
@@ 1. Introduction
Recommender systems are pervasive across digital platforms, tailoring ranked lists of items for users in diverse contexts such as Spotify, YouTube, TikTok, Facebook, and Amazon. Notably, point-of-interest (POI) recommender systems suggest locations such as restaurants and stores, prioritizing geographical proximity (Ye et al., 2011). These systems significantly influence user experience and decision-making, necessitating an emphasis on fairness to avoid perpetuating biases. The discourse on fairness within machine learning and its applications to recommender systems has expanded, addressing different fairness definitions and sources of bias (Ekstrand et al., 2022; Mehrabi et al., 2021).
The advancement of multi-stakeholder recommendation systems introduces the challenge of balancing diverging interests between consumers and providers. While consumers benefit from recommendations that accurately reflect their preferences and intentions, providers gain from equitable exposure and opportunities presented by their items being recommended. This dynamic creates a necessity for systems that can manage and mitigate the conflicting needs of all parties involved (Burke et al., 2018; Abdollahpouri et al., 2020). Our approach integrates a ’fairness factor’ into POI recommender systems, applying it through post-processing or post-filtering techniques that remain agnostic to the underlying model, thus ensuring broad applicability.
This paper focuses on exploring multi-objective fairness in POI recommender systems, with particular attention to the balance between provider and consumer interests. We pose the following research questions to guide our investigation:
  • How does the scoring model used for the provider fairness factor influence the overall performance of the recommender system?
  • What are the effects of implementing a provider fairness factor on performance, fairness, and distance metrics, and how do these compare when a consumer fairness factor is applied?
  • How does the recommender system adapt when both provider and consumer fairness factors are integrated simultaneously?
@@ 2. Modelling the Task and the Fairness Factors
This project aims to create a recommender system model that incorporates a user $u\in U$, a point-of-interest (POI) $p\in P$, and contextual factors $C(u,p)$ (e.g., distance, time, social connections). Additionally, it integrates consumer and provider fairness via abstract functions $F_{c}(u)$ for consumer fairness and $F_{p}(u)$ for provider fairness. This model, denoted $\hat{M}$, integrates the aforementioned contexts to produce personalized POI recommendations. Fairness is woven into this framework by adjusting the ranking score as follows:
@@
The tunable parameters $\alpha$ and $\beta$ determine the impact of fairness on the rankings. Different configurations of these parameters will be evaluated to illustrate their effects on system performance and fairness, aiming to identify optimal balances in a Pareto front.
In this front, we propose a framework CAPRI-FAIR [^1] https://github.com/cruiseresearchgroup/CAPRI-FAIR/, where we extend a pre-existing framework for context-aware POI recommendation and incorporate fairness factors in a similar fashion. Consequently, we also add evaluations that take into account user and item fairness.
@@
Each POI has an associated popularity, indicated by user check-ins, which follows a power law distribution. Most check-ins are concentrated among a few POIs, while a large majority fall into the ”long tail” with minimal exposure. This characteristic of the distribution underscores the challenge of achieving fairness among providers, as most POIs receive scant attention. To address this, we model popularity with a power law, where the score of a POI is inversely proportional to its check-in count, helping to elevate less popular POIs. For a given number of $x$ check-ins, the estimated number of POIs $y$ with that popularity is expressed inline as $y=w_{0}\cdot x^{w_{1}}$. This relationship undergoes a log-log transformation to fit a linear regression, where parameters are determined using L2 regularization to manage data skew, with a regularization factor of 10.0 ensuring stable predictions. In addition to the power law model, we explore linear and logistic regression models for comparison. The linear model does not scale down exposure as sharply with increasing popularity, potentially offering a more gradual adjustment for moderately popular POIs. In contrast, logistic regression treats slight increases in popularity more harshly, impacting even moderately popular POIs significantly.
To assess provider fairness, we measure each item’s exposure across recommendations, assuming each POI represents a unique provider. Exposure is influenced by the ranking position, which can be modeled as a function. For simplicity it is binary in the paper, indicating the presence or absence in the recommendation list.
@@ 3. Experiment Setup
Dataset Model Exposure Model Precision Long-tail Exposure
0 0.25 0.50 0.75 1 0 0.25 0.50 0.75 1
Yelp GeoSoCa Linear 0.0225 0.0134 0.0117 0.0111 0.0107 1.7010 3.4467 4.1525 4.5026 4.6888
Logistic 0.0225 0.0173 0.0089 0.0073 0.0073 1.7010 3.0696 4.5693 4.6969 4.6970
PowerLaw 0.0225 0.0137 0.0109 0.0096 0.0081 1.7010 3.1636 3.5278 2.5025 1.8429
LORE Linear 0.0230 0.0179 0.0147 0.0132 0.0124 1.5290 2.7975 3.6141 4.1361 4.4426
Logistic 0.0230 0.0183 0.0096 0.0074 0.0072 1.5290 3.4217 4.8180 4.9503 4.9565
PowerLaw 0.0230 0.0196 0.0136 0.0106 0.0082 1.5290 2.2284 2.1647 1.8887 1.5768
USG Linear 0.0297 0.0314 0.0278 0.0228 0.0186 0.0077 0.0341 0.5424 1.8214 3.0765
Logistic 0.0297 0.0296 0.0257 0.0123 0.0042 0.0077 0.0461 1.1532 4.5227 5.6810
PowerLaw 0.0297 0.0300 0.0291 0.0152 0.0036 0.0077 0.0209 0.2791 2.9418 3.9288
Gowalla GeoSoCa Linear 0.0381 0.0300 0.0287 0.0274 0.0266 0.7990 0.9796 1.0734 1.1626 1.2465
Logistic 0.0381 0.0316 0.0182 0.0140 0.0132 0.7990 1.2142 1.9263 2.0336 2.0459
PowerLaw 0.0381 0.0277 0.0242 0.0222 0.0203 0.7990 1.3231 1.8073 2.0207 2.0636
LORE Linear 0.0482 0.0427 0.0405 0.0389 0.0378 0.7354 0.9356 1.0757 1.1958 1.2961
Logistic 0.0482 0.0283 0.0149 0.0127 0.0125 0.7354 1.6326 2.0425 2.0815 2.0842
PowerLaw 0.0482 0.0392 0.0260 0.0191 0.0156 0.7354 1.2983 1.6527 1.7724 1.7771
USG Linear 0.0514 0.0504 0.0477 0.0446 0.0424 0.0766 0.1056 0.1567 0.2523 0.3689
Logistic 0.0514 0.0500 0.0334 0.0195 0.0103 0.0766 0.2319 1.4315 2.0190 2.2654
PowerLaw 0.0514 0.0525 0.0468 0.0293 0.0158 0.0766 0.1513 0.7876 1.6531 2.0789
Table 1. Precision and Long-tail Exposure for different Provider Fairness factor weights and Exposure models used in the CAPRI-FAIR framework.
@@
We employ the Gowalla(Cho et al., 2011) and Yelp(Liu et al., 2017) datasets, recognized for their comprehensive feature sets which facilitate model comparisons (Werneck et al., 2021; Rahmani et al., 2022a). Following standard preprocessing, we remove POIs visited by fewer than 10 users and users visiting fewer than 10 POIs. The data is split into training, validation, and testing sets using a 70-20-10 ratio. Users are categorized as ‘active’ or “inactive’ based on their activity, with the top 20% being active. This is based on the defaults given by the original CAPRI framework (Tourani et al., 2023), as well as the Pareto principle (i.e. where 20% of users account for 80% of check-ins). Items are similarly divided into ‘short-head’ and ‘long-tail’ groups, based on the top 20% of popularity metrics.
@@
We measure both Precision and Exposure across different groups to evaluate user and item fairness. We employ the Generalized Cross-Entropy (GCE), using Pearson’s $\chi^{2}$ measure for its robustness to outliers (Rahmani et al., 2022b; Deldjoo et al., 2021), to assess fairness. This measure helps quantify disparities in item exposure, particularly focusing on long-tail exposure improvements. For consumer fairness, recommendation gain for each user is defined based on the presence of recommended POIs in their previous visits, akin to an attention function used for item exposure. Similarly, provider fairness is evaluated by whether POIs appear in recommendation lists, indicating their exposure. These metrics are crucial for understanding the impact of fairness adjustments on the recommendation system.
@@
Our model is compared against three commonly used POI recommenders: USG (Ye et al., 2011), GeoSoCa(Zhang and Chow, 2015), and LORE(Zhang et al., 2014), which are widely used in recent studies (Tourani et al., 2023; Werneck et al., 2021; Yu et al., 2020; Rahmani et al., 2022c). In order to maintain consistency, we utilize the same datasets as referenced in (Tourani et al., 2023). We investigate several key hypotheses on the impact of fairness factors on model performance and fairness metrics:
  • Exposure Model Impact: We evaluate three exposure models—power law, linear, and logistic—across values of $\alpha$ (0, 0.25, 0.5, 1.0) to determine their effect on precision and long-tail exposure at $k=10$. Statistical tests ascertain if the choice of exposure model significantly influences these metrics.
  • Provider Fairness Factor: This factor’s impact on long-tail item exposure is assessed over a range of weights from 0 to 1, analyzing whether improvements in exposure can be achieved without compromising accuracy.
  • Consumer Fairness Factor: We examine how adjusting the consumer fairness factor improves precision for inactive users, varying the weight from 0 to 1 and monitoring impacts at $k=10$.
Each hypothesis is rigorously tested to find an optimal balance between fairness and performance, ensuring the model’s practical applicability and theoretical soundness.
@@ 4. Results and Discussion
Table 2. Precision, GCE, and Mean median distance metrics with respect to different weights for providers and consumers. Yelp dataset.
The benchmark models were run first without modification, then slowly increasing the provider fairness factor’s coefficient $\alpha$ as outlined above. The impact of the provider fairness model used is first assessed by taking note of its impact on the precision and exposure metrics for the long tail, which it is aiming to improve. The results are shown below in Table 1.
The score of the best-performing exposure model for the metric is highlighted in bold. For precision, which is the accuracy or performance metric of choice here, the linear exposure model performs the best in most cases. This makes sense as it doesn’t punish even slightly popular POIs as harshly as the other models, as shown by the curves in Figure 5 in the appendix. Likewise, the logistic model performs best regarding exposure to the long tail, which makes sense as it only provides weights to the least popular POIs. However, note that the linear model still reasonably increases the long-tail exposure as the weight increases. The power law model, however, seems to perform the worst for the GeoSoCa and LORE models and even decreases past a certain weight. Running a Kruskal-Wallis test on the precision and exposure values reveals a statistically significant difference in metric values among the 3 exposure models, starting from a factor weight of $\alpha=0.5$, indicating that the choice of the exposure model does make a difference in the precision and long-tail exposure metrics. We can now use the linear model for the provider fairness factor, as it offers the highest precision values and permits moderate long-tail exposure.
@@
After adopting the linear model as the provider fairness factor for further analysis, its effects on performance metrics were thoroughly evaluated. As the previous table shows, increasing $a$ typically leads to a decline in precision. However, we see that it does help with exposure to the long-tail POIs, who, under the normal regime, experience very low exposure rates of 1-2 recommendation appearances overall. For USG, this exposure is as low as near-0 by default. As shown in Figure 1, an increase in the factor’s weight increases the long-tail exposure, reaching up to 1.4 times the default for GeoSoCa and LORE and an even larger increase for the USG model.
@@
Increasing the weight of the consumer fairness factor, denoted as $\beta$, primarily aims to enhance precision for inactive users without affecting active users, as shown in Figures 2. In the Yelp dataset, adjusting $\beta$ initially raises precision for the GeoSoCa and LORE models from 2.3% to 2.6%, but further increases lead to performance degradation, especially in the USG model, where precision drops from 3.0% to 2.8%. A Wilcoxon signed rank test confirms that even a slight increase in $\beta$ causes significant changes in precision across models, highlighting the sensitivity of recommender systems to fairness adjustments.
@@
Naturally, the impact of combining provider and consumer fairness factors on performance metrics can be explored through GCE metrics for different tradeoffs, which are illustrated in Figure 3 and detailed in Table 2. The figure effectively plots the tradeoff between user precision GCE and item exposure GCE metrics, using colour coding to denote precision levels and various shapes to differentiate between models and $k$ values. This visualization is particularly revealing across the Yelp and Gowalla datasets, where the Pareto front usually occupies the upper right region of each plot. Here, models that achieve high user GCE often do so at the expense of item GCE, indicating a significant tradeoff. This tradeoff is particularly stark for the USG model, which generally shows the best precision, as evidenced by its performance metrics detailed in Table 2. The table highlights that GeoSoCa and LORE perform best with full weight on consumer fairness, but this undermines provider fairness. In contrast, the USG model works better with equal weights on both fairness factors. Adjusting fairness weights has nuanced impacts on mean median distances, typically improving provider fairness across all models. This illustrates how each model’s properties, combined with dataset characteristics, affect the balance between consumer and provider fairness strategies, impacting overall fairness and performance in POI recommendation systems.

### v2

@@ Abstract.
Point-of-interest (POI) recommendation considers spatio-temporal constraints like distance, peak hours, and user check-ins. Given their influence on both consumer experience and POI business, it’s crucial to consider fairness from multiple perspectives. Unfortunately, these systems often provide less accurate recommendations to inactive users and less exposure to unpopular POIs. This paper develops a post-filter methodology that integrates provider and consumer fairness into existing models, aiming to balance fairness metrics like item exposure with performance metrics such as precision and distance. Experiments show that a linear scoring model for provider fairness in re-scoring items offers the best tradeoff between performance and long-tail exposure, sometimes without significant precision loss. Addressing consumer fairness by recommending more popular POIs to inactive users increased precision in some models and datasets. However, combinations that reached the Pareto front of consumer and provider fairness yielded the lowest precision values, underscoring that tradeoffs depend heavily on the model and dataset.
@@
journalyear: 2024copyright: acmlicensedconference: 18th ACM Conference on Recommender Systems; October 14–18, 2024; Bari, Italybooktitle: 18th ACM Conference on Recommender Systems (RecSys ’24), October 14–18, 2024, Bari, Italydoi: 10.1145/3640457.3688170isbn: 979-8-4007-0505-2/24/10ccs: Information systems Recommender systemsccs: Information systems Combination, fusion and federated search
@@ 1. Introduction
Recommender systems are pivotal in digital environments, tailoring content across platforms like Spotify, YouTube, TikTok, and Facebook. Specifically, point-of-interest (POI) recommender systems suggest locations such as restaurants and stores, emphasizing geographical proximity as a crucial factor (Ye et al., 2011). While these systems significantly enhance user experience, the critical issue of fairness within these algorithms has gained substantial attention. Research in machine learning fairness has evolved, addressing various definitions and the multifaceted nature of biases in recommender systems (Ekstrand et al., 2022; Mehrabi et al., 2021). Existing work (Burke et al., 2018; Abdollahpouri et al., 2020) has begun to address multi-stakeholder recommendation frameworks that consider the needs of both consumers and providers. Consumers benefit most from recommendations that accurately reflect their interests and intentions, while providers gain from equitable exposure of their items. Despite these advancements, current models often struggle to balance these interests effectively without sacrificing the quality of recommendations or exacerbating existing biases.
Our research targets this gap by integrating a ‘fairness factor’ into POI recommender systems through post-processing techniques adaptable to various underlying models. This approach aims to refine the balance between provider and consumer interests, often overlooked in conventional systems. Additionally, we explore multi-objective fairness in POI recommender systems. This area has seen limited investigation, particularly in how different fairness objectives can be simultaneously optimized without detriment to the overall system performance (Patro et al., 2020a; Naghiaei et al., 2022).
This paper contributes to the field by proposing a methodology that reconciles the often conflicting objectives of different stakeholders within the POI recommendation context. By doing so, we aim to push the boundaries of what is currently possible in fairness-oriented recommender systems, providing a template for future research to build upon (Patro et al., 2020b; Merinov, 2023). We pose the following research questions to guide our investigation:
  • RQ1: How does the scoring model for the provider fairness factor affect the recommender system’s performance?
  • RQ2: What are the impacts of the provider fairness factor on performance, fairness, and distance metrics compared to the consumer fairness factor?
  • RQ3: How does the recommender system perform when integrating provider and consumer fairness factors?
Aligned with these questions, our investigation will test several hypotheses:
  • H1: The exposure model has an impact on the system’s performance.
  • H2: The provider fairness factor improves long-tail item exposure and the consumer fairness factor enhances precision for inactive users, both without compromising accuracy.
  • H3: Integrating both provider and consumer fairness factors potentially improves GCE fairness metrics.
@@ 2. Modelling the Task and the Fairness Factors
This project aims to develop a recommender system model that incorporates a user $u\in U$, a point-of-interest (POI) $p\in P$, and contextual factors $C(u,p)$ (e.g., distance, time, social connections), all of which are normalized to ensure values range between 0 and 1. Additionally, it integrates consumer and provider fairness through abstract functions $F_{c}(u)$ and $F_{p}(u)$, which also range from 0 to 1 for consumer and provider fairness, respectively. The model, denoted by $\hat{M}$, combines these contexts to produce personalized POI recommendations. Fairness is incorporated into this framework by adjusting the ranking score as follows:
@@
where the tunable parameters $\alpha$ and $\beta$ modulate the impact of fairness considerations and ensure the score $\mathbf{M}(u,p)$ remains within the range of 0 to 1. Different configurations will be evaluated to illustrate their effects on system performance and fairness, aiming to identify optimal balances in a Pareto front.
In this front, we propose a framework CAPRI-FAIR [^1] https://github.com/cruiseresearchgroup/CAPRI-FAIR, where we extend a pre-existing framework for context-aware POI recommendation and incorporate fairness factors. We also add evaluations that consider user and item fairness.
@@
Each POI has an associated popularity, indicated by user check-ins, which follows a power law distribution. Most check-ins are concentrated among a few POIs, while a large majority fall into the ‘long tail’ with minimal exposure. This distribution characteristic underscores the challenge of achieving fairness among providers, as most POIs receive scant attention. To address this, we model popularity with a power law, where the score of a POI is inversely proportional to its check-in count, helping to elevate less popular POIs. For a given number of $x$ check-ins, the estimated number of POIs $y$ with that popularity is expressed inline as $y=w_{0}\cdot x^{w_{1}}$. This relationship undergoes a log-log transformation to fit a linear regression, where parameters are determined using L2 regularization to manage data skew, with a regularization factor of 10.0 ensuring stable predictions. In addition to the power law model, we explore linear and logistic regression models for comparison. The linear model does not scale down exposure as sharply with increasing popularity, offering a more gradual adjustment for moderately popular POIs. In contrast, logistic regression treats slight increases in popularity more harshly, impacting even moderately popular POIs significantly. To assess provider fairness, we measure each item’s exposure across recommendations, where each POI is assumed to represent a unique provider. Exposure ($E$) is quantified using an attention function $a(p,P_{R})$, which, for simplicity, is binary in our experiments, indicating whether an item $p$ is present or absent in the recommendation list. The exposure of an item $p$ is defined as $(E_{p}=\sum_{P_{R}\subset P}a(p,P_{R}))$, aligning with methods in related work that scale exposure based on item visibility or rank within the list, facilitating comparison of different ranking algorithms’ impact on provider fairness.
@@ 3. Experiment Setup
We employ the Gowalla(Cho et al., 2011) and Yelp(Liu et al., 2017) datasets, recognized for their comprehensive feature sets which facilitate model comparisons (Werneck et al., 2021; Rahmani et al., 2022a). Following standard preprocessing, we remove POIs visited by fewer than 10 users and users who have visited fewer than 10 POIs. After this filtering process, the dataset includes 7135 users for Yelp and 5628 users for Gowall. We divide each user’s sequence of POI visits into 70% training, 20% validation, and 10% testing, maintaining the chronological order to simulate real-world scenarios and ensure robust evaluations of our recommendation system. Users are categorized as ‘active’ or “inactive’ based on their activity, with the top 20% being active. This is based on the defaults given by the original CAPRI framework (Tourani et al., 2023) and the Pareto principle (i.e., where 20% of users account for 80% of check-ins). Items are similarly divided into ‘short-head’ and ‘long-tail’ groups based on the top 20% of popularity metrics.
Table 1. Precision and Long-tail Exposure by Provider Fairness Weights and Exposure Models in CAPRI-FAIR framework.
Dataset Yelp Gowalla
Model Exposure Model Precision Long-tail Exposure Precision Long-tail Exposure
0.25 0.50 0.75 1 0.25 0.50 0.75 1 0.25 0.50 0.75 1 0.25 0.50 0.75 1
GeoSoCa Linear 0.0134 0.0117 0.0111 0.0107 3.4467 4.1525 4.5026 4.6888 0.0300 0.0287 0.0274 0.0266 0.9796 1.0734 1.1626 1.2465
Logistic 0.0173 0.0089 0.0073 0.0073 3.0696 4.5693 4.6969 4.6970 0.0316 0.0182 0.0140 0.0132 1.2142 1.9263 2.0336 2.0459
PowerLaw 0.0137 0.0109 0.0096 0.0081 3.1636 3.5278 2.5025 1.8429 0.0277 0.0242 0.0222 0.0203 1.3231 1.8073 2.0207 2.0636
LORE Linear 0.0179 0.0147 0.0132 0.0124 2.7975 3.6141 4.1361 4.4426 0.0427 0.0405 0.0389 0.0378 0.9356 1.0757 1.1958 1.2961
Logistic 0.0183 0.0096 0.0074 0.0072 3.4217 4.8180 4.9503 4.9565 0.0283 0.0149 0.0127 0.0125 1.6326 2.0425 2.0815 2.0842
PowerLaw 0.0196 0.0136 0.0106 0.0082 2.2284 2.1647 1.8887 1.5768 0.0392 0.0260 0.0191 0.0156 1.2983 1.6527 1.7724 1.7771
USG Linear 0.0314 0.0278 0.0228 0.0186 0.0341 0.5424 1.8214 3.0765 0.0504 0.0477 0.0446 0.0424 0.1056 0.1567 0.2523 0.3689
Logistic 0.0296 0.0257 0.0123 0.0042 0.0461 1.1532 4.5227 5.6810 0.0500 0.0334 0.0195 0.0103 0.2319 1.4315 2.0190 2.2654
PowerLaw 0.0300 0.0291 0.0152 0.0036 0.0209 0.2791 2.9418 3.9288 0.0525 0.0468 0.0293 0.0158 0.1513 0.7876 1.6531 2.0789
@@
We measure both Precision and Exposure across different groups to evaluate user and item fairness. For fairness assessment, we employ the Generalized Cross-Entropy (GCE) with Pearson’s $\chi^{2}$ measure, noted for its robustness to outliers (Rahmani et al., 2022b; Deldjoo et al., 2021). This measure quantifies item exposure disparities, focusing on long-tail exposure improvements. For consumer fairness, recommendation gain for each user is defined based on the presence of recommended POIs in their previous visits. Similarly, provider fairness is evaluated by whether POIs appear in recommendation lists. Each POI’s latitude and longitude are used to compute distances, and the user’s position is averaged from the coordinates of their previously visited POIs. Distances are then calculated using the standard geographical distance formula, ensuring accurate measurement of proximity.
@@
Our model is benchmarked against three widely used POI recommenders (Tourani et al., 2023; Werneck et al., 2021; Yu et al., 2020; Rahmani et al., 2022c): USG (Ye et al., 2011), GeoSoCa (Zhang and Chow, 2015), and LORE (Zhang et al., 2014). USG integrates user preferences, social influence, and geographic data. GeoSoCa leverages geographical, social, and categorical correlations, while LORE combines geographic, temporal, and social contexts. These methods are chosen for their robust integration of diverse contextual factors crucial for effective POI recommendations. Future evaluations may extend to include graph-based, deep learning-based, and generative AI-based methods for a more comprehensive comparison. Consistent datasets are used as per (Tourani et al., 2023) to maintain comparability.
Our experiments systematically evaluate the effects of fairness factors on model performance and fairness metrics. We assess three exposure models—power law, linear, and logistic—across $\alpha$ values from 0 to 1, analyzing their impact on precision and long-tail exposure at $k=10$. Both provider and consumer fairness factors are explored for their influence on item exposure and precision for inactive users. We employ the Kruskal-Wallis test to assess differences among the three exposure models starting at $\alpha=0.25$. Pairwise Mann-Whitney U tests further analyze specific model comparisons, enhancing our evaluation of each model’s impact.
@@ 4. Results and Discussion
Table 2. Precision, GCE, and Mean Median Distance Metrics for Different Weights of Providers and Consumers, Yelp Dataset.
@@
The benchmark models were initially run without modifications, then progressively with increased provider fairness factor $\alpha$ coefficients to assess the impact delineated in H1. The first assessment focused on the effect of the provider fairness model on precision and long-tail exposure metrics, which it aims to improve. Results presented in Table 1 highlight the outcomes of these evaluations.
For precision, the key performance metric, the linear exposure model consistently outperforms others, as indicated in bold in the results. This model’s relative leniency on moderately popular POIs is illustrated by the curves in Figure 5 in the appendix. Conversely, the logistic model boosts exposure for the least popular POIs. The power law model, however, underperforms for the GeoSoCa and LORE models, particularly at higher $\alpha$ weights. Differences between the exposure models validate H1 that the choice of exposure model is crucial. Specifically, the precision of the linear model is statistically significantly greater than that of the power law model. Pairwise comparisons between the linear and logistic models and logistic and power law models show no significant differences. Therefore, the linear model is chosen for the provider fairness factor due to its superior performance. We acknowledge this choice’s limitations and will include a comprehensive table of all pairwise results across models in Appendix A.3 for detailed comparison.
@@
After adopting the linear model as the provider fairness factor, we observed its effects on performance metrics related to H2. As shown in Table 1, increasing $\alpha$ reduces precision but significantly enhances long-tail exposure. For instance, in the USG model, where baseline exposure is nearly zero, the increase in $\alpha$ substantially boosts exposure, as detailed in Figure 1. This supports H2, demonstrating that provider fairness can enhance visibility for underrepresented POIs without overly compromising precision.
@@
Increasing the weight of the consumer fairness factor, denoted as $\beta$, aims to enhance precision for inactive users, aligning with H3. As depicted in Figure 2, adjusting $\beta$ in the Yelp dataset initially improves precision for the GeoSoCa and LORE models from 2.3% to 2.6%. However, further increases cause precision to decline, notably in the USG model, where it drops from 3.0% to 2.8%. A Wilcoxon signed rank test confirms significant changes in precision across models with minor adjustments in $\beta$, underscoring the impact of fairness modifications on recommender systems.
@@
Exploring the impact of combining provider and consumer fairness factors, as outlined in H3, can be seen through GCE metrics for different tradeoffs, detailed in Figure 3 and Table 2. This figure plots the tradeoff between user precision GCE and item exposure GCE metrics, using color to denote precision levels and shapes to differentiate between models and $k$ values. Across the Yelp and Gowalla datasets, the Pareto front typically occupies the upper right region of each plot, highlighting a significant tradeoff where high user GCE often comes at the expense of item GCE, especially evident in the USG model which shows optimal precision. Table 2 further reveals that while GeoSoCa and LORE achieve the best precision with full weight on consumer fairness, this approach diminishes provider fairness. Conversely, the USG model demonstrates better performance with balanced fairness weights. Such adjustments to fairness weights reveal nuanced impacts on mean median distances, generally enhancing provider fairness across all models. This analysis underscores how specific properties of each model and dataset characteristics influence the effectiveness of combined fairness strategies, thus affecting overall outcomes in POI recommendation systems. For more details, see Appendix A.4.